"Cosy dy\/dx" = "Sin^2x Cosx"


1
Expert's answer
2021-05-14T11:50:15-0400

Given,

"cosy\cdot \dfrac{dy}{dx}=sin^2x\times cosx\\\ \\cosy\times dy=sin^2x\times cosxdx"


Now, Integrating both sides

"\int cosydy=\int sin^2xcosxdx"

"Let \ I_1=\int sin^2xcosxdx\\"

Substitute sinx = t

then cosxdx = dt

"I_1=\int t^2dt\\I_1=\frac{t^3}{3}\\"


Now, "I_1= \dfrac{1}{3}sin^3x"


So,

"\int cosydy=\int sin^2xcosxdx"

"siny =\dfrac{1}{3}sin^3x +C\\\boxed{y=sin^{-1}(\frac{1}{3}sin^3x+C)}"